Find an orthogonal matrix associated with matrix A D

1 2
2 4 
:
1
Expert's answer
2020-04-09T14:47:50-0400

Find an orthogonal matrix B such that "P^TAP" is diagonal when


"A=\begin{bmatrix}
 1 & 2 \\
 2 & 4
\end{bmatrix}"

First, the find eigenvalues and eigenvectors

Start from forming a new matrix by subtracting "\lambda" from the diagonal entries of the given matrix:


"A-\lambda I=\begin{bmatrix}
 1-\lambda & 2 \\
 2 & 4-\lambda
\end{bmatrix}"

Find the determinant of the obtained matrix:


"det(A-\lambda I)=\begin{vmatrix}
 1-\lambda & 2 \\
 2 & 4-\lambda
\end{vmatrix}=""=(1-\lambda)(4-\lambda)-2(2)=\lambda^2-5\lambda"

This is a characteristic polynomial.

Solve the equation


"\lambda^2-5\lambda=0"

The roots are:

"\lambda_1=5"

"\lambda_2=0"

These are the eigenvalues.

Next, find the eigenvectors.

"\lambda_1=5"


"\begin{bmatrix}
 1-\lambda & 2 \\
 2 & 4-\lambda
\end{bmatrix}=\begin{bmatrix}
 -4 & 2 \\
 2 & -1
\end{bmatrix}"

Perform row operations to obtain the rref of the matrix:

"R_2=R_2+\dfrac{1}{2}R_1:"


"\begin{bmatrix}
 -4 & 2 \\
 0 & 0
\end{bmatrix}"

"R_2=-\dfrac{1}{4}R_1:"


"\begin{bmatrix}
 1& - {1 \over 2}\\
 0 & 0
\end{bmatrix}"

Now, solve the matrix equation


"\begin{bmatrix}
 1& - {1 \over 2}\\
 0 & 0
\end{bmatrix}\begin{bmatrix}
 v_1\\
 v_2
\end{bmatrix}=\begin{bmatrix}
 0\\
 0
\end{bmatrix}"

If we take "v_2=t," then "v_1={1 \over2}t, v_2=t"

Therefore


"\text{v}=\begin{bmatrix}
 t/2\\
 t
\end{bmatrix}=\begin{bmatrix}
 1/2\\
 1
\end{bmatrix}t"

"\lambda_2=0"


"\begin{bmatrix}
 1-\lambda & 2 \\
 2 & 4-\lambda
\end{bmatrix}=\begin{bmatrix}
 1 & 2 \\
 2 & 4
\end{bmatrix}"

Perform row operations to obtain the rref of the matrix:

"R_2=R_2-(2)R_1:"


"\begin{bmatrix}
 1 & 2 \\
 0 & 0
\end{bmatrix}"

Now, solve the matrix equation


"\begin{bmatrix}
 1& 2\\
 0 & 0
\end{bmatrix}\begin{bmatrix}
 v_1\\
 v_2
\end{bmatrix}=\begin{bmatrix}
 0\\
 0
\end{bmatrix}"

If we take "v_2=t," then "v_1=-2t, v_2=t".

Therefore


"\text{v}=\begin{bmatrix}
-2t\\
 t
\end{bmatrix}=\begin{bmatrix}
 -2\\
 1
\end{bmatrix}t"

Eigenvalue: 5, eigenvector: "\begin{bmatrix}
 1/2\\
 1
\end{bmatrix}"

Eigenvalue: 0, eigenvector:"\begin{bmatrix}
 -2\\
 1
\end{bmatrix}"

Form the matrix "P," whose i-th column is the i-th eigenvector:


"\begin{bmatrix}
 1/2& -2\\
 1 & 1
\end{bmatrix}"

"(1\/2)^2+(1)^2=5\/4"

"(-2)^2+(1)2=5"

Then an orthogonal matrix "P"


"P=\begin{bmatrix}
 1/\sqrt{5} & -2/\sqrt{5}\\
 2/\sqrt{5} & 1/\sqrt{5}
\end{bmatrix}"